OpenStudy (anonymous):

A pick-up truck is coasting at a speed Va along a straight and level road. When a load equivalent to 10% of the truck's mass is thrown off the bed, parallel to the ground and in the forward direction, the truck is brought to a halt. if the direction in which this mass is thrown is exactly reversed, but the speed of this mass relative to the truck remains the same, the wagon accelerates to a new speed Vb. Calculate the ratio Vb/Va.

OpenStudy (anonymous):

Vb = 100/81Va ? Since kinetic energy = (m*v^2)/2 (m1*v1^2)/2=(m2*v2^2)/2